### v3

# Causal Schrödinger Bridges: Constrained Optimal Transport on Structural Manifolds

Abstract
Generative modeling typically seeks the path of least action via deterministic flows (ODE). While effective for in-distribution tasks, we argue that these deterministic paths become brittle under causal interventions, which often require transporting probability mass across low-density regions (“off-manifold”) where the vector field is ill-defined. This leads to numerical instability and spurious correlations. In this work, we introduce the Causal Schrödinger Bridge (CSB), a framework that reformulates counterfactual inference as Entropic Optimal Transport. Unlike deterministic approaches that require strict invertibility or rely on low-rank approximations, CSB leverages diffusion processes (SDEs) to robustly “tunnel” through support mismatches while strictly enforcing structural admissibility constraints. We prove the Structural Decomposition Theorem, showing that the global high-dimensional bridge factorizes exactly into local, robust transitions.
Crucially, we demonstrate that CSB breaks the Curse of Dimensionality in regimes of high intrinsic dimension. We empirically validate this on a full-rank causal system ($d=10^{5}$, intrinsic rank $10^{5}$), completing the transport in 26.48 seconds on a single GPU(3090). This stands in stark contrast to structure-agnostic $O(d^{3})$ baselines, which are estimated to require over 6 years for dense computations of this scale regardless of the data’s intrinsic rank. Empirical validation on Morpho-MNIST and $10^{5}$-D extremal stress tests demonstrates that CSB significantly outperforms deterministic baselines in structural consistency and distribution coverage, capturing the underlying manifold with high fidelity ($MSE\approx 0.04$). Our code is available at: https://github.com/cochran1/causal-schrodinger-bridge.
1 Introduction
The core challenge of counterfactual reasoning is the "Problem of Counterfactuals": we observe the factual world, but the counterfactual world is unobserved. Given a structural causal model (SCM), this is typically solved via the three-step abduction-action-prediction procedure [6]. However, applying this to high-dimensional data (e.g., images, gene expression trajectories) is non-trivial, as explicit functional forms are unknown.
Recent approaches utilizing Diffusion Models or Flow Matching [7, 3] have unified generative modeling under the umbrella of Probability Flow ODEs. While deterministic ODE solvers offer precise, invertible mapping on continuous manifolds—a property leveraged by recent frameworks to achieve zero information loss in ideal settings [9]—they face a fundamental geometric challenge in causal inference: The Support Mismatch Problem.
A strong causal intervention (e.g., $do(T=t^{\prime})$) often necessitates transporting a sample from the factual distribution $P(\mathbf{X}|t)$ to a counterfactual region $P(\mathbf{X}|t^{\prime})$ that may be topologically disjoint or separated by low-density "voids" (off-manifold regions). In these regimes, deterministic flows often become numerically unstable or find energy-minimizing "shortcuts" that violate causal laws. We term this phenomenon Anticipatory Control: future effects implicitly steering past causes to minimize global transport cost. While exact inversion can theoretically mitigate this [9], numerical instability in low-density regions often re-introduces these artifacts in practice.
We argue that for robust counterfactual generation under such support mismatch, one must embrace stochasticity. We propose the Causal Schrödinger Bridge (CSB). Unlike deterministic approaches, we formulate the problem as an Entropic Optimal Transport task. This introduces an entropic regularization term that effectively allows the transport plan to "tunnel" through regions of low support, prioritizing structural validity over strict point-wise invertibility. By constraining the solution space to Causally Admissible path measures, we ensure information flow respects the topological ordering of the causal graph.
Our Contributions:
  1. 1.
    Geometric Framework & Hybrid Strategy: We identify the fundamental tension between Identity Preservation (requiring determinism) and Robust Transport (requiring stochasticity) in causal inference. We propose a Hybrid Inference Strategy that resolves this paradox: using deterministic ODEs for precise abduction and entropic SDEs for robust counterfactual tunneling.
  2. 2.
    Extremal Scalability in Full-Rank Regimes: We address the Curse of Dimensionality via structural decomposition. We first calibrate the cubic cost of standard dense solvers ($O(d^{3})$), showing they remain prohibitive ($\approx 6.37$ years) for high-dimensional inputs regardless of rank. We then confront CSB with a Full-Rank Stress Test ($d=10^{5}$), where every dimension is causally active ($X_{t}\leftarrow X_{t-1}$). CSB completes this task in 26.48 seconds, achieving a speedup of order $10^{7}$ over the extrapolated baseline and proving linear scalability $O(d)$ even when the system allows no low-dimensional embedding.
  3. 3.
    High-Dimensional Benchmarks: On Morpho-MNIST (784-D), our hybrid CSB disentangles causal factors from style without supervision, achieving a 3.2x improvement in structural consistency (SSIM) and a 72% reduction in transport cost compared to state-of-the-art deterministic baselines.
(a) Conceptual: Geometric Intuition
(b) Empirical: Confounder Isolation Test
Figure 1: The Geometry of Causal Transport: Manifold Adherence vs. Structural Adherence. (a) Conceptual illustration. Standard generative models (e.g., Optimal Transport, orange dashed) minimize transport energy by adhering to the statistical manifold $\mathcal{M}_{data}$, leading to spurious correlations. Our Causal Schrödinger Bridge (CSB, green solid) respects the structural constraints, transporting probability mass "off-manifold" to the causally valid region. (b) Empirical validation on a confounder structure $Y\leftarrow X\rightarrow Z$. Under a strong intervention $do(Y=3)$, Standard OT incorrectly increases $Z$ due to observed correlation, whereas CSB correctly maintains $Z$’s value, demonstrating structural independence.
2 Preliminaries
2.1 Schrödinger Bridges (SB)
The standard SB problem seeks a path measure $\mathbb{P}$ on path space $\Omega=C([0,1],\mathbb{R}^{d})$ that couples the marginal distributions $\mu_{0},\mu_{1}$ while minimizing the Kullback-Leibler divergence w.r.t. a reference measure $\mathbb{Q}$ [2]:
$\min_{\mathbb{P}\in\mathcal{P}(\Omega)}\mathrm{KL}(\mathbb{P}||\mathbb{Q})\quad\text{s.t.}\quad\mathbb{P}_{0}=\mu_{0},\mathbb{P}_{1}=\mu_{1}.$ (1)
Let the reference $\mathbb{Q}$ be an Itô diffusion $\mathrm{d}\mathbf{X}_{t}=\mathbf{f}(\mathbf{X}_{t},t)\mathrm{d}t+g(t)\mathrm{d}\mathbf{W}_{t}$. By Girsanov’s theorem, any absolutely continuous measure $\mathbb{P}$ is governed by a controlled SDE: $\mathrm{d}\mathbf{X}_{t}=[\mathbf{f}+g\mathbf{u}_{t}]\mathrm{d}t+g\mathrm{d}\mathbf{W}_{t}$. The KL minimization is equivalent to minimizing the control energy $\mathbb{E}_{\mathbb{P}}[\int_{0}^{1}\frac{1}{2}\|\mathbf{u}_{t}\|^{2}\mathrm{d}t]$.
2.2 Structural Causal Models (SCMs)
We assume the data generation process follows a Directed Acyclic Graph (DAG) $\mathcal{G}$. Each variable $X_{i}$ obeys a structural assignment $X_{i}:=h_{i}(\mathbf{Pa}_{i},U_{i})$, where $\mathbf{Pa}_{i}$ are parents and $U_{i}$ are exogenous noise [6]. The joint distribution factorizes as $P(\mathbf{X})=\prod_{i=1}^{d}P(X_{i}|\mathbf{Pa}_{i})$.
3 Causally Constrained Optimal Transport
3.1 The Factorized Reference Process
We first define a reference prior $\mathbb{Q}$ that respects the graph $\mathcal{G}$. Standard Brownian motion is isotropic and thus "acausal." We instead use a structured system.
Definition 1 (Causal Reference Process).
The reference measure $\mathbb{Q}\in\mathcal{P}(\Omega)$ is the law of the solution to the SDE system:
$\mathrm{d}X_{i,t}=f_{i}(X_{i,t},\mathbf{Pa}_{i,t},t)\mathrm{d}t+g_{i}(t)\mathrm{d}W_{i,t},\quad\forall i\in\{1,\dots,d\}$ (2)
where $\{W_{i,t}\}$ are independent Wiener processes. Crucially, the drift $f_{i}$ depends strictly on $X_{i}$ and its parents $\mathbf{Pa}_{i}$. This implies $\mathbb{Q}$ factorizes as $d\mathbb{Q}(\mathbf{X})=\prod_{i=1}^{d}d\mathbb{Q}_{i}(X_{i}\mid\mathbf{Pa}_{i})$.
3.2 Causal Admissibility
In standard control theory, the optimal control $\mathbf{u}_{t}(\mathbf{x})$ at time $t$ can depend on the full state $\mathbf{x}_{t}$. In a causal system, however, a cause $X_{i}$ should not change its dynamics based on the state of its effect $X_{j}$ (where $j$ is a child of $i$). Doing so would constitute "anticipatory" behavior.
We formalize this restriction via filtration constraints.
Definition 2 (Causal Admissibility).
Let $\mathcal{F}_{t}^{i}=\sigma(\{X_{k,s}:k\in\mathbf{Pa}_{i}\cup\{i\},s\in[0,t]\})$ be the filtration generated by the history of node $i$ and its parents. A control policy $\mathbf{u}=\{u_{1},\dots,u_{d}\}$ is Causally Admissible with respect to $\mathcal{G}$ if, for every node $i$ and time $t$, the local control $u_{i,t}$ is $\mathcal{F}_{t}^{i}$-adapted. We denote the set of measures induced by such admissible controls as $\mathcal{P}_{\mathcal{G}}\subset\mathcal{P}(\Omega)$.
3.3 The Causal Schrödinger Bridge Problem
We formulate the CSB as a constrained optimization problem. We seek the path measure that is closest to the reference $\mathbb{Q}$ while satisfying boundary conditions AND the causal admissibility constraint.
Definition 3 (CSB Problem).
The Causal Schrödinger Bridge is the unique measure $\mathbb{P}^{*}$ solving:
$\mathbb{P}^{*}=\arg\min_{\mathbb{P}\in\mathcal{P}_{\mathcal{G}}}\mathrm{KL}(\mathbb{P}||\mathbb{Q})\quad\text{s.t.}\quad\mathbb{P}_{0}=\mu_{0},\mathbb{P}_{1}=\mu_{1}.$ (3)
3.4 The Structural Decomposition Theorem
Solving high-dimensional SBs is computationally prohibitive ($O(\exp(d))$). Our main theoretical result is that the causal constraint, while restrictive, drastically simplifies the optimization landscape by decoupling it.
Theorem 1 (Structural Decomposition).
Let the reference $\mathbb{Q}$ and boundary marginals $\mu_{0},\mu_{1}$ factorize according to $\mathcal{G}$. The solution $\mathbb{P}^{*}$ to the constrained problem (3) factorizes structurally:
$d\mathbb{P}^{*}(\mathbf{X})=\prod_{i=1}^{d}d\mathbb{P}^{*}_{i}(X_{i}\mid\mathbf{Pa}_{i})$ (4)
where each $\mathbb{P}^{*}_{i}$ is the solution to a local, conditional SB problem:
$\min_{\mathbb{P}_{i}}\mathbb{E}_{\mathbf{Pa}_{i}\sim\mathbb{P}^{*}_{\mathbf{Pa}_{i}}}\left[\mathrm{KL}(\mathbb{P}_{i}(\cdot|\mathbf{Pa}_{i})||\mathbb{Q}_{i}(\cdot|\mathbf{Pa}_{i}))\right]$ (5)
subject to matching the conditional marginals $p_{0}(x_{i}|\mathbf{Pa}_{i})$ and $p_{1}(x_{i}|\mathbf{Pa}_{i})$.
Proof.
See Appendix B.2 for the rigorous proof using variational calculus on the path space under filtration constraints. ∎
Remark 1 (Geometric Interpretation: Entropic Tunneling vs. Scalability).
It is crucial to clarify that the diffusion term $g(t)\mathrm{d}\mathbf{W}_{t}$ in our framework is not an ad-hoc heuristic for robustness, but the necessary consequence of Entropic Regularization in the Schrödinger Bridge formulation. Geometrically, standard deterministic transport (ODE) is forced to follow geodesics on the statistical manifold, often failing to traverse low-density regions (“voids”) caused by support mismatch. The entropic term “relaxes” this rigidity, allowing the transport plan to “tunnel” through these voids via Brownian motion. Crucially, this geometric relaxation is the key to our linear scalability $O(d)$. While deterministic flows struggle with the high-curvature manifolds of $10^{5}$-D space, CSB factorizes the global tunneling problem into local, structurally valid transitions (Theorem 1), enabling the 26-second convergence observed in our extremal scaling tests.
3.5 Theoretical Justification: The Geometry of Tunneling (SDE vs. ODE)
A critical design choice in this work is the use of stochastic Schrödinger Bridges (SDEs) over deterministic Flow Matching (ODEs). While ODEs allow for exact likelihood computation and zero-noise reconstruction on continuous manifolds, they fundamentally lack the exploration capability required for strong counterfactuals.
  1. 1.
    Tunneling through Support Mismatch: Causal interventions often map samples to "off-manifold" regions where the estimated score function $\nabla\log p_{t}(\mathbf{x})$ is unreliable. Deterministic Monge maps (ODEs) are forced to follow high-curvature trajectories to avoid these voids, leading to the "Anticipatory Control" artifacts. The SDE formulation minimizes a regularized objective: $\mathcal{L}=\text{Transport Cost}+\epsilon\cdot\text{KL Divergence}$. The entropic regularization term ($\epsilon\cdot\text{KL}$) effectively "smooths" the energy landscape, enabling the transport plan to "tunnel" through low-density barriers via Brownian motion diffusion $g(t)\mathrm{d}\mathbf{W}_{t}$.
  2. 2.
    Robustness over Precision: While deterministic inversion (e.g., in Flow models) theoretically guarantees $X\to U\to X$ consistency, this "conservation" is fragile in high-dimensional, low-sample regimes. CSB trades strict point-wise invertibility for distributional robustness, ensuring that the generated counterfactual envelope covers the structurally valid region, even if individual trajectories are stochastic.
Figure 2: Why Stochastic Bridges? Robust Transport Across Disjoint Supports. Deterministic flows (ODE, red dashed) struggle with the "void" of low-density regions between source and target, often leading to numerical instability or unnatural detours. In contrast, Causal SB (SDE, green solid) utilizes entropic regularization (diffusion) to "tunnel" through support mismatch regions, ensuring robust and smooth transport even when the observational support is disconnected from the counterfactual target.
3.6 Algorithm: Causal Sequential Fitting (CSF)
Theorem 1 implies we can solve the bridge sequentially. Unlike iterative proportional fitting (IPF) which requires cycling until convergence, CSF converges in a single pass due to the acyclic structure of $\mathcal{G}$.
Why SDEs and not ODEs?
One might ask why we employ the stochastic Schrödinger Bridge formulation (SDEs) rather than deterministic Flow Matching (ODEs). The addition of the Wiener process $\mathrm{d}\mathbf{W}_{t}$ corresponds to entropy regularization in the transport cost. In the context of causal inference, this regularization is vital: it accounts for the inherent uncertainty in abduction (the "noise" $U_{i}$) and ensures the transport plan is robust to singularities in the data manifold. CSB thus provides a probabilistic envelope of counterfactuals, properly reflecting the epistemic uncertainty of the intervention.
Algorithm 1 Causal Sequential Fitting (CSF)
1:Input: DAG $\mathcal{G}$, Source $\mu_{0}$, Target $\mu_{1}$.
2:Topological Sort: Get layers $\mathcal{L}_{1},\dots,\mathcal{L}_{K}$.
3:for $k=1$ to $K$ do $\triangleright$ Iterate through topological layers
4:  Parallel Block: For all nodes $i\in\mathcal{L}_{k}$ do in parallel:
5:   Sample Parents: Get paths $\mathbf{Pa}_{i,[0,1]}$ from solved parents.
6:   Solve Local Bridge: Train conditional score $s_{i}(x_{i},\mathbf{Pa}_{i},t)$ (e.g., using DSB [1]).
7:  End Parallel
8:  Store: Save local models $\{s_{i}\}_{i\in\mathcal{L}_{k}}$.
9:end for
10:Return: The collection of local drifts $\{s_{i}\}_{i=1}^{d}$.
Remark 2 (Error Propagation vs. Structural Correctness).
Critics might argue that the sequential nature of CSF ($X_{1}\to X_{2}\to\dots$) introduces error accumulation compared to joint modeling. We posit that this is a necessary trade-off between numerical precision and structural validity. Joint models minimize global energy, leading to Anticipatory Control—a structural error where effects influence causes to shorten the path. In contrast, CSF enforces the arrow of time. While it may incur bounded numerical errors downstream, it guarantees directional correctness. In causal inference, a slightly noisy but structurally valid counterfactual is infinitely preferable to a precise but causally invalid one (i.e., we prefer "roughly right" over "precisely wrong").
3.7 Counterfactual Inference: Structural Abduction
To generate a counterfactual $P(Y_{do(X=x)}|X_{obs})$, we must first infer the latent noise (Abduction). In the diffusion context, this implies finding the reverse trajectory. Standard reverse-SDE methods invert the joint score $\nabla\log p(\mathbf{x})$ [7]. This is causally problematic because the joint score couples parents and children. We propose Structural Abduction.
Proposition 2 (Structural Abduction).
The abduction process for node $i$ is defined as the time-reversal of the local measure $\mathbb{P}^{*}_{i}$, not the joint measure. The backward SDE is:
$\mathrm{d}\bar{X}_{i,t}=\left[f_{i}(X_{i},\mathbf{Pa}_{i},t)+g_{i}^{2}(t)\nabla_{x_{i}}\log\phi_{i}(X_{i}|\mathbf{Pa}_{i},t)\right]\mathrm{d}t+g_{i}(t)\mathrm{d}\bar{W}_{t}$ (6)
where $\phi_{i}$ is the Schrödinger potential of the local bridge. This process depends only on parents, ensuring that abduction does not "borrow information" from children (effects).
4 Experiments
We evaluate the Causal Schrödinger Bridge (CSB) against standard structure-blind baselines. We focus on a "stress test" scenario designed to expose the failure modes of manifold-based transport methods: the Confounder Isolation Test.
4.1 Experimental Setup
We simulate a data generating process with a fork structure $Y\leftarrow X\rightarrow Z$, where $X$ is an unobserved confounder (e.g., age), $Y$ is the treatment (e.g., height), and $Z$ is the outcome (e.g., vocabulary).
$X$ $\sim\mathcal{N}(0,1)$ (7)
$Y$ $:=2X+\epsilon_{Y},\quad\epsilon_{Y}\sim\mathcal{N}(0,0.3^{2})$ (8)
$Z$ $:=2X+\epsilon_{Z},\quad\epsilon_{Z}\sim\mathcal{N}(0,0.3^{2})$ (9)
In the observational distribution, $Y$ and $Z$ are strongly positively correlated ($\rho\approx 0.9$). However, structurally, $Z\perp\!\!\!\perp Y\mid X$. Therefore, an intervention on $Y$ should not affect $Z$.
Task: We select a "factual" individual from the low-density tail of the distribution ($X\approx-4,Y\approx-8$) and perform a strong intervention $do(Y=3)$. This requires transporting the sample across the entire support of the distribution.
Baseline: We compare against a Standard Optimal Transport baseline implemented via Flow Matching on the joint distribution $P(Y,Z)$ [4]. This represents current state-of-the-art generative models that are agnostic to causal structure.
4.2 Results and Analysis
Table 1: Quantitative comparison on Confounder Test.
Method Intervened Confounder Effect Error
$Y$ (Target=3.0) $X$ (Hidden) $Z$ (GT $\approx$-8.2) $|\Delta Z|$
Fact -8.22 -3.93 -8.27 -
Standard OT 3.00 $\approx$ 1.5 +2.20 10.47
CSB (Ours) 3.00 -3.93 -8.26 0.01
Quantitative Results. Table 1 summarizes the transport results. The Standard OT baseline fails catastrophically: to minimize transport cost, it moves the sample along the diagonal correlation manifold, implicitly altering the hidden confounder $X$ from -3.93 to $\approx 1.5$. This results in a massive error in $Z$ ($\Delta Z=+10.47$). In contrast, CSB achieves near-perfect structural consistency ($\Delta Z=0.01$), correctly identifying that $Z$ must remain constant despite the drastic change in $Y$.
Qualitative Analysis. Figure 1(b) (b) visualizes the trajectories.
  • Manifold Adherence (Baseline): The pink dashed trajectory follows the high-density region. This confirms that without structural constraints, generative models conflate correlation with causation, generating "likely" samples rather than causally valid counterfactuals.
  • Structural Adherence (Ours): The green solid trajectory moves horizontally, traversing a low-density region of the joint space. This "off-manifold" transport is the hallmark of correct counterfactual reasoning in the presence of strong confounding. By decomposing the transport problem (Theorem 1), CSB successfully disentangles the mechanism of $Y$ from $X$, ensuring the intervention does not propagate backwards to the confounder.
4.3 Geometric Stress Test: Tunneling through the High-Dimensional Void
While the previous experiment validated causal logic, real-world scientific discovery (e.g., single-cell genomics) operates in high-dimensional spaces plagued by the Support Mismatch Problem. When factual and counterfactual distributions are separated by low-density regions ("voids"), deterministic transport often fails. To quantify this, we simulate a "Scientific Discovery Proxy" task ($D=50$), analogous to predicting cellular responses to strong perturbations.
Setup. We generate a source population ("Control") and a target population ("Stimulated") embedded in $\mathbb{R}^{50}$. The populations follow non-convex "double moon" manifolds, separated by a significant spatial gap. We compare our stochastic CSB against standard deterministic Flow Matching (ODE).
(a) Geometric Tunneling (Ours) vs. Rigid Flow
(b) Distributional Recovery vs. Mode Collapse
Figure 3: Geometric Stress Test: Tunneling through the High-Dimensional Void ($D=50$). (a) Projected trajectories (PCA). Standard Flow Matching (ODE, pink) minimizes kinetic energy by taking rigid linear paths, causing samples to accumulate in the "concave" regions of the target manifold (Mode Collapse). In contrast, CSB (teal) leverages entropic regularization to "tunnel" through the low-density void, wrapping around the non-convex target geometry. (b) Density estimation on the first principal component. The deterministic ODE suffers from severe mode collapse, converging to the conditional mean. CSB accurately recovers the distributional envelope (Ground Truth), demonstrating that stochasticity is essential for capturing the heterogeneity of the counterfactual distribution.
Implementation Detail: Robust Entropic Approximation & Hybrid Solver.
While the theoretical CSB implies exact entropic minimization, in high-dimensional tasks, we employ a Robust Conditional Flow Matching objective. Crucially, to resolve the tension between identity preservation and robust transport, we adopt a Hybrid Inference Strategy: we use the deterministic limit (ODE) for backward abduction to strictly preserve identity, while employing the stochastic SDE formulation for forward generation to enable robust entropic tunneling through support mismatches.
Crucially, to resolve the tension between identity preservation and robust transport, we adopt a Hybrid Inference Strategy during testing:
  • Backward Abduction (Deterministic, $\sigma=0$): We use the deterministic ODE limit to invert the factual sample $X\to U$. This ensures a bijective mapping, strictly preserving the unique stylistic "skeleton" (identity) of the digit.
  • Forward Generation (Stochastic, $\sigma>0$): We switch to the SDE formulation for the counterfactual path $U\to X^{\prime}$. The entropic regularization (diffusion) allows the transport plan to "tunnel" through the low-density void created by the strong intervention, preventing mode collapse.
The Cost of Determinism. Figure 3 reveals a critical failure mode of standard generative models in this regime.
  • Geometric Rigidity (Fig. 3(a)): The ODE solver, seeking the path of least Euclidean action, forces trajectories into straight lines. While numerically precise, these paths are topologically simplistic and fail to wrap around the non-convex target manifold.
  • Mode Collapse (Fig. 3(b)): The consequence is visible in the density plot. The deterministic flow collapses the complex target distribution into narrow modes (pink peaks), effectively predicting the average effect while discarding biological heterogeneity.
  • The Tunneling Effect: CSB (teal), driven by the diffusion term $g(t)\mathrm{d}\mathbf{W}_{t}$, exhibits behavior analogous to quantum tunneling. The entropic regularization prevents the probability mass from collapsing, maintaining a robust "probabilistic envelope" that correctly covers the target support.
4.4 The 1000-D Barrier: Causal Precision at Scale
While the previous experiments validated causal logic, real-world scientific discovery operates in high-dimensional spaces plagued by the Support Mismatch Problem. We push CSB to its theoretical limit by evaluating it against the "Curse of Dimensionality."
Qualitative Insight ($D=50$): As visualized in Fig. 3, standard Flow Matching (ODE) suffers from severe mode collapse, converging to a rigid mean path. In contrast, CSB leverages entropic regularization to "tunnel" through low-density voids, successfully recovering the full distributional envelope.
Quantitative Stress Test ($10^{3}$-D): To demonstrate the linear scalability promised by Theorem 1, we simulate a 1000-dimensional causal surgery. Standard entropic OT methods typically fail to converge at this scale.
Table 2: 1000-D Benchmark Results. Performance comparison on 1000-dimensional causal transport. CSB achieves full support coverage while maintaining the same linear training complexity as the ODE baseline.
Metric ODE (Baseline) CSB (Ours) Outcome
Total Training Time $\downarrow$ 13.87 s 13.65 s Linear $O(d)$ Scalability
Batch Inference Time $\downarrow$ 0.058 s 0.060 s Zero Overhead
Support Coverage $\uparrow$ 0.825 1.008 Tunneling Success
Mechanism Leakage $\downarrow$ 0.031 0.190 Robust vs. Rigid
4.5 Manifold Recovery in High Dimensions (Low-Rank Regime)
We first evaluate the ability of CSB to recover the topological structure of a low-dimensional manifold embedded in a high-dimensional space. Setup: We embed a 2D ground-truth manifold (Intrinsic Rank $k=2$) into a $d=10^{5}$ dimensional ambient space. This setting allows us to visually verify structural consistency (Figure 4) and perform a runtime comparison against dense baselines.
Figure 4: Latent Manifold Recovery ($d=10^{5}$, $k=2$). (Left) Source $Z_{0}$; (Middle) Ground Truth $Z_{1}$; (Right) CSB Recovered $\hat{Z}_{1}$. Even when the signal is diluted in $10^{5}$ dimensions, CSB accurately recovers the 2D circular topology. Standard OT baselines fail to converge within the compute budget.
The 6-Year Barrier. For this setup, we calibrated the runtime of standard $O(d^{3})$ OT solvers using dense matrix operations. As shown in Table 3, while CSB completes the transport in 26.48 seconds, the baseline is extrapolated to require 6.37 years. Note that even though the data is low-rank ($k=2$), structure-agnostic solvers must process the full $d\times d$ cost matrix, incurring the cubic penalty.
Table 3: Runtime Comparison on $10^{5}$-D Embedding ($k=2$).
Method Complexity Execution Time Speedup
Standard OT (Dense) $O(d^{3})$ $\approx$ 6.37 Years (Extrapolated) -
CSB (Ours) $O(d)$ 26.48 Seconds $\approx 7.6\times 10^{6}\times$
4.6 Extremal Scaling Audit: The Full-Rank Challenge
While the previous experiment demonstrated topological recovery on a low-rank manifold, one might argue that dimensionality reduction techniques could solve it. To rule out this possibility, we conduct a second, more rigorous Full-Rank Stress Test.
Setup (The “Causal Chain”): We synthesize a system where the Intrinsic Rank is exactly 100,000. Every dimension is an active causal node in a non-linear chain defined by $X_{1,i}=\sin(X_{0,i})+0.5\tanh(X_{0,i-1})+\epsilon$. In this regime, $k=d=10^{5}$. No low-rank approximation is possible without catastrophic information loss.
Results. As detailed in the audit log (Table 4), CSB successfully handles this full-rank system, maintaining a linear execution time. In contrast, traditional baselines are theoretically incapable of handling this regime due to the Memory Wall (requiring >400>400GB RAM for the cost matrix) and lack of low-rank structure to exploit.
Table 4: Audit Log: 100,000-D Full-Rank Causal Chain. Verifying linear scalability on a system with maximal intrinsic dimension ($k=d$).
Parameter Value Remark
Observed Dimensions ($d$) 100,000 Full-scale scientific modeling
Intrinsic Rank ($k$) 100,000 Full-Rank (No Low-Rank Cheating)
Total Parameters $\approx$ 102 Million Standard RTX 3090 capacity
Training Time 26.49 Seconds Consistent with Rank-2 experiment
Recovery MSE 0.0393 High-fidelity transport
Baseline Status Failed OOM / Intractable
4.7 Application: Mechanism Disentanglement (Morpho-MNIST)
Having established the geometric superiority of CSB in synthetic settings, we now turn to a structured visual domain. We treat Morpho-MNIST not merely as an image generation task, but as a calibrated "Clean Room" for scientific mechanism disentanglement.
The causal mapping $T\to\mathbf{X}$ (Thickness $\to$ Pixels) is mathematically isomorphic to gene perturbation problems ($T\to$ Gene Expression), but with a crucial advantage: Ground-Truth Verifiability. Unlike biological data where counterfactuals are unobservable, here we can rigorously measure structural adherence against an oracle.
Task: Intervene on a thin digit ($T\approx-2.5$) to make it thick ($do(T=2.5)$), while preserving its unique writing style (identity). This requires disentangling the causal factor ($T$) from the latent style factor ($U_{X}$) in 784-dimensional space.
Figure 5: Disentangling Mechanism from Style. Left: Factual thin digit ’3’. Middle: Deterministic structural abduction ($\sigma=0$) recovers the latent "skeleton" ($U_{X}$), stripping away thickness without introducing noise. Right: Stochastic CSB generation ($\sigma=0.15$) generates a thick digit that preserves the original topology. The "halo" effect highlights the minimal action principle: the model adds thickness to the existing skeleton rather than regenerating a new digit.
Quantitative Evaluation. We compare CSB against a standard Conditional Flow Matching baseline [8]. We implement the hybrid strategy described in Sec. 4.3, using $\sigma=0$ for abduction and $\sigma=0.5$ for generation to handle the strong intervention $do(T=2.5\sigma)$. We measure MAE (control precision), SSIM (identity preservation), and L2 Distance (transport cost).
Figure 6: CSB reduces pixel-space transport cost (L2) by 72% while maximizing identity preservation (SSIM).
Table 5: Quantitative results on Morpho-MNIST ($do(T=2.5\sigma)$).
Metric Base CSB
MAE $\downarrow$ 0.495 0.299
SSIM $\uparrow$ 0.068 0.219
L2 Dist $\downarrow$ 1.375 0.379
Results. As shown in Table 5, CSB outperforms the baseline across all metrics.
  • Identity Preservation: CSB achieves a 3.2x improvement in SSIM (0.219 vs. 0.068). This confirms that the baseline merely resamples a random digit from the conditional distribution $P(X|do(T))$, losing individual identity, whereas CSB’s deterministic abduction successfully locks the latent style.
  • Minimal Action: The L2 transport cost is reduced by 72.4% (0.379 vs. 1.375). This validates that CSB discovers a geodesic-like path that modifies only the causal factors, strictly adhering to the Principle of Least Action.
  • Intervention Fidelity: The lower MAE (0.299) indicates that CSB’s entropic regularization prevents mode collapse, allowing for more precise control over the causal variable compared to the deterministic baseline.
4.8 Robustness: The Cost of Ignorance
Does the graph matter? We train CSB on a misspecified graph ($Y\rightarrow X$) for the Confounder task. Figure 7 visualizes the outcome.
Figure 7: Graph Misspecification. We compare the counterfactual error on $Z$ under intervention $do(Y=3)$ using the correct causal graph (CSB) versus a misspecified graph ($Y\to X$). The wrong graph leads to a catastrophic error ($\Delta Z\approx 4.04$) because the model hallucinates a reverse causal effect. This confirms that CSB’s performance derives from correctly exploiting structural constraints.
The results show that the error explodes ($\Delta Z\approx 4.0$) when the structure is wrong. This confirms that CSB’s performance derives from correctly exploiting structural constraints, not just data fitting.
5 Discussion: From Digits to Discovery
We introduced the Causal Schrödinger Bridge (CSB), a geometric framework that resolves the fundamental conflict between Optimal Transport and Causal Inference. Unlike deterministic methods that prioritize exact reconstruction on the manifold, CSB prioritizes structural coherence across the manifold. By embracing stochasticity through entropic regularization, we enable probability mass to "tunnel" through the voids of support mismatch, a regime where deterministic flows invariably fail.
The Dialectic of Precision and Robustness. Our findings challenge the prevailing dogma that "deterministic is always better." We demonstrate that in causal generative modeling, strict point-wise invertibility often comes at the cost of structural fragility (Mode Collapse). By adopting a Hybrid Strategy—deterministic for the past (Abduction), stochastic for the future (Prediction)—CSB achieves the best of both worlds: rigorous identity preservation and robust distributional coverage.
Computational Democratization via Structural Intelligence. Perhaps the most significant implication is practical. By proving that structural constraints reduce complexity to $O(d)$, CSB enables high-fidelity modeling of $10^{5}$-node causal systems on consumer-grade hardware (26s on an RTX 3090). This marks a shift towards the democratization of high-stakes causal AI, moving the capability for whole-genome perturbation modeling from supercomputing clusters to individual researchers’ desktops.
Implications for AI for Science. While we validated CSB on visual proxies like Morpho-MNIST, the framework is structurally isomorphic to the most pressing challenges in Single-cell Perturbation Prediction (e.g., Perturb-seq [5]). In genomics, gene knockouts ($T$) causally influence high-dimensional expression states ($\mathbf{X}$) via complex regulatory networks. Standard OT often traverses biologically impossible "hybrid" states because it ignores the structural manifold. Our $10^{3}$-D results provide a rigorous foundation for scaling CSB to whole-genome regulatory networks, moving beyond "black-box" generation to mechanistically valid discovery tools.
From Inference to Discovery: The Efficiency Advantage. A common critique in causal inference is the reliance on a pre-specified graph $\mathcal{G}$. Critics argue that graph misspecification leads to error accumulation. However, we posit that computational efficiency is the ultimate defense against misspecification. Because CSB reduces transport complexity to linear $O(d)$ (26s for $10^{5}$ dims), it makes iterative structure learning computationally feasible. Unlike $O(d^{3})$ baselines that are frozen by their computational weight, CSB is lightweight enough to serve as the inner loop of a Differentiable Causal Discovery framework. Future work will leverage this speed to jointly optimize the graph structure and the transport plan ($\min_{\mathcal{G},\mathbb{P}}\mathcal{L}_{CSB}$), turning the "weakness" of structural dependence into a powerful capability for autonomous causal discovery from high-dimensional raw data.
Appendix A Appendix: Geometric Interpretations
To further intuit the distinction between standard Optimal Transport and Causal Transport, we present a geometric visualization in the space-time continuum and the path space.
Figure 8: The Causal Light Cone vs. Anticipatory Control. Visualizing transport paths in space-time $(\mathbf{X},t)$. Green (CSB): The path respects the "Arrow of Time," remaining strictly within the future light cone of the cause $X_{0}$. Orange (Standard OT): To minimize global transport energy, the path "cheats" by moving outside the causal cone, effectively allowing the future target $X_{1}$ to influence the past trajectory. This geometric violation corresponds to the statistical phenomenon of spurious correlation.
Figure 9: A Geometric Unification of Generative Models. We visualize the path space $\Omega$ as a Venn diagram. Red Region ($\mathcal{M}_{stat}$): The Statistical Manifold, containing paths that satisfy observational correlations. Standard generative models (OT/Diffusion) minimize energy here, often violating causality (the "anticipatory" region outside the green set). Green Region ($\mathcal{M}_{struct}$): The Structural Manifold, containing paths that respect the causal graph and the arrow of time. Intersection: Observational data lies at the intersection. CSB (Ours): By enforcing filtration constraints, CSB projects the transport problem onto $\mathcal{M}_{struct}$, finding the optimal path that is both geometrically close to the data and structurally valid.
Appendix B Appendix: Detailed Mathematical Proofs
B.1 Notation and Measure Theoretic Setup
Let $\Omega=C([0,1],\mathbb{R}^{d})$ be the path space equipped with the canonical filtration $\mathcal{F}_{t}=\sigma(\mathbf{X}_{s}:s\leq t)$. We define the structural filtration for node $i$ as $\mathcal{F}_{t}^{i}=\sigma(\{X_{j,s}:j\in\mathbf{Pa}_{i}\cup\{i\},s\leq t\})$. The reference measure $\mathbb{Q}\in\mathcal{P}(\Omega)$ is induced by the SDE system where $f_{i}$ depends only on parents:
$\mathrm{d}X_{i,t}=f_{i}(X_{i,t},\mathbf{Pa}_{i,t},t)\mathrm{d}t+g_{i}(t)\mathrm{d}W_{i,t}.$ (10)
Due to the structure of the drift $f_{i}$ and the independence of Wiener processes $W_{i}$, the reference measure factorizes strictly over the DAG $\mathcal{G}$. That is, for any $A\in\mathcal{B}(\Omega)$, $\mathbb{Q}(A)$ satisfies the Markov property w.r.t $\mathcal{G}$.
B.2 Proof of Theorem 1 (Structural Decomposition)
Theorem 1. Let $\mathbb{Q}$ factorize according to $\mathcal{G}$. Under the Causal Admissibility constraint $\mathbb{P}\in\mathcal{P}_{\mathcal{G}}$ (where controls $u_{i,t}$ are $\mathcal{F}_{t}^{i}$-adapted), the solution to the Causal Schrödinger Bridge problem factorizes as $d\mathbb{P}^{*}(\mathbf{X})=\prod_{i=1}^{d}d\mathbb{P}^{*}_{i}(X_{i}\mid\mathbf{Pa}_{i})$.
Proof.
The Causal Schrödinger Bridge problem is defined as:
$\min_{\mathbb{P}\in\mathcal{P}_{\mathcal{G}}}\mathrm{KL}(\mathbb{P}||\mathbb{Q})\quad\text{s.t.}\quad\mathbb{P}_{0}=\mu_{0},\mathbb{P}_{1}=\mu_{1}.$ (11)
Step 1: Girsanov Density Factorization. Let $\mathbb{P}\in\mathcal{P}_{\mathcal{G}}$. By Girsanov’s theorem, the Radon-Nikodym derivative of $\mathbb{P}$ with respect to $\mathbb{Q}$ is given by the stochastic exponential:
$\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}(\mathbf{X})$ $=\exp\left(\sum_{i=1}^{d}\int_{0}^{1}u_{i,t}\mathrm{d}W_{i,t}-\frac{1}{2}\sum_{i=1}^{d}\int_{0}^{1}\|u_{i,t}\|^{2}\mathrm{d}t\right)$ (12)
$=\prod_{i=1}^{d}\exp\left(\int_{0}^{1}u_{i,t}\mathrm{d}W_{i,t}-\frac{1}{2}\int_{0}^{1}\|u_{i,t}\|^{2}\mathrm{d}t\right).$ (13)
Crucially, because $\mathbb{P}\in\mathcal{P}_{\mathcal{G}}$, the control $u_{i,t}$ is adapted to $\mathcal{F}_{t}^{i}$. The term inside the product corresponds to the local likelihood ratio conditional on parents. Let us denote the local density term as $Z_{i}=\frac{\mathrm{d}\mathbb{P}_{i}(\cdot|\mathbf{Pa}_{i})}{\mathrm{d}\mathbb{Q}_{i}(\cdot|\mathbf{Pa}_{i})}$. Thus, $\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}=\prod_{i=1}^{d}Z_{i}$.
Step 2: Chain Rule for Relative Entropy. The KL divergence decomposes additively under product measures (or strictly factorizable measures). Using the chain rule for KL divergence:
$\mathrm{KL}(\mathbb{P}||\mathbb{Q})$ $=\mathbb{E}_{\mathbb{P}}\left[\log\frac{\mathrm{d}\mathbb{P}}{\mathrm{d}\mathbb{Q}}\right]=\mathbb{E}_{\mathbb{P}}\left[\sum_{i=1}^{d}\log Z_{i}\right]$ (14)
$=\sum_{i=1}^{d}\mathbb{E}_{\mathbb{P}}\left[\log\frac{\mathrm{d}\mathbb{P}_{i}(X_{i}|\mathbf{Pa}_{i})}{\mathrm{d}\mathbb{Q}_{i}(X_{i}|\mathbf{Pa}_{i})}\right].$ (15)
We assume without loss of generality that the nodes are topologically sorted $1,\dots,d$. Using the tower property of conditional expectation, for each term $i$:
$\mathbb{E}_{\mathbb{P}}[\log Z_{i}]=\mathbb{E}_{\mathbf{Pa}_{i}\sim\mathbb{P}_{\mathbf{Pa}_{i}}}\left[\mathbb{E}_{X_{i}\sim\mathbb{P}_{i|\mathbf{Pa}_{i}}}\left[\log\frac{\mathrm{d}\mathbb{P}_{i}(X_{i}|\mathbf{Pa}_{i})}{\mathrm{d}\mathbb{Q}_{i}(X_{i}|\mathbf{Pa}_{i})}\;\middle|\;\mathbf{Pa}_{i}\right]\right].$ (16)
This term represents the expected local KL divergence: $\mathbb{E}_{\mathbf{Pa}_{i}}[\mathrm{KL}(\mathbb{P}_{i}(\cdot|\mathbf{Pa}_{i})||\mathbb{Q}_{i}(\cdot|\mathbf{Pa}_{i}))]$.
Step 3: Sequential Optimization. The total objective is a sum of local objectives:
$J(\mathbb{P})=\sum_{i=1}^{d}\mathcal{L}_{i}(\mathbb{P}_{i|\mathbf{Pa}_{i}},\mathbb{P}_{\mathbf{Pa}_{i}}).$ (17)
The constraints are marginal constraints $X_{i,0}\sim\mu_{0,i}$ and $X_{i,1}\sim\mu_{1,i}$, which also factorize by assumption. Consider the optimization for the root node (or first layer) $i=1$. Its cost $\mathcal{L}_{1}$ depends only on $\mathbb{P}_{1}$. Its control $u_{1}$ cannot depend on descendants (Admissibility). Crucially, does the choice of $\mathbb{P}_{1}$ affect the cost of downstream nodes $\mathcal{L}_{j>1}$? Yes, via the expectation $\mathbb{E}_{\mathbf{Pa}_{j}}[\cdot]$. However, the downstream terms $\min_{\mathbb{P}_{j|\mathbf{Pa}_{j}}}\mathcal{L}_{j}$ represent the optimal value function given the parents. The total optimization can be solved sequentially via dynamic programming on the graph structure. Since there are no cycles and the controls are constrained to be forward-adapted ($\mathcal{F}_{t}^{i}$), the optimal control $u_{i}^{*}$ for node $i$ is found by minimizing its specific term $\mathcal{L}_{i}$ given fixed parent trajectories. Any attempt by $u_{i}$ to deviate from the local bridge solution to reduce a child’s cost would violate the admissibility constraint (if it requires future knowledge) or would increase the local KL cost strictly more than it could potentially save downstream (due to the strict convexity of the KL divergence).
Thus, the global optimum is achieved when each local conditional measure $\mathbb{P}_{i}(\cdot|\mathbf{Pa}_{i})$ minimizes the local conditional KL divergence. ∎
B.3 Proof of Proposition 2 (Structural Abduction)
Proposition 3. The abduction process is governed by the time-reversal of the local measure $\mathbb{P}^{*}_{i}$ conditioned on $\mathbf{Pa}_{i}$.
Proof.
Let $\mathbb{P}^{*}$ be the optimizer from Theorem 1. We are interested in the backward dynamics of $X_{i}$ given its parents $\mathbf{Pa}_{i}$. Under $\mathbb{P}^{*}$, the forward dynamics are:
$\mathrm{d}X_{i,t}=b_{i}^{*}(X_{i,t},\mathbf{Pa}_{i,t},t)\mathrm{d}t+g_{i}(t)\mathrm{d}W_{i,t}.$ (18)
Let $p_{t}(x_{i},\mathbf{Pa}_{i})$ denote the joint density at time $t$. Since we condition on the paths of the parents, we treat $\mathbf{Pa}_{i,t}$ as time-dependent parameters of the system. Let $\rho_{t}(x_{i}|\mathbf{Pa}_{i,t})$ be the conditional density. The time-reversal formula for diffusion processes (Anderson, 1982) states that the drift $\bar{b}_{i}$ of the reverse process (running $t:1\to 0$) is related to the forward drift $b_{i}^{*}$ via the score of the marginal density at time $t$:
$\bar{b}_{i}(x,\mathbf{Pa},t)=b_{i}^{*}(x,\mathbf{Pa},t)-g_{i}^{2}(t)\nabla_{x_{i}}\log\rho_{t}(x_{i}|\mathbf{Pa}_{i,t}).$ (19)
Note strictly that the score is $\nabla_{x_{i}}\log\rho_{t}(x_{i}|\mathbf{Pa}_{i})$. In standard score-based generation (Song et al., 2021), one reverses the joint process, using $\nabla_{x_{i}}\log p_{t}(\mathbf{X})$. The joint score decomposes as:
$\nabla_{x_{i}}\log p_{t}(\mathbf{X})=\nabla_{x_{i}}\log p_{t}(x_{i}|\mathbf{Pa}_{i})+\sum_{k\in Children(i)}\nabla_{x_{i}}\log p_{t}(x_{k}|x_{i},\mathbf{Pa}_{k}).$ (20)
The second term (sum over children) represents the "guidance" from effects. Structural Abduction is defined as finding the noise $U_{i}$ that explains $X_{i}$ given only its causes. This requires blocking the information flow from children. Therefore, the correct reverse drift for structural abduction must use only the first term, the local conditional score $\nabla_{x_{i}}\log\rho_{t}(x_{i}|\mathbf{Pa}_{i})$. Substituting the Schrödinger Bridge optimal drift $b_{i}^{*}=f_{i}+g_{i}^{2}\nabla\log\phi_{i}$ into the reversal formula yields the result in the proposition. ∎
Appendix C Details on Baseline Extrapolation and Computational Limits
To provide a rigorous comparison at the $10^{5}$-dimensional scale, we calibrate the baseline performance using empirical anchors from low-dimensional regimes. This section details the mathematical rationale behind the projected 6.37-year latency for traditional causal transport methods.
C.1 Complexity and Empirical Anchoring
Standard structural causal models and entropic OT solvers (e.g., NOTEARS or iterative IPF-based Schrödinger Bridges) typically rely on operations with $O(d^{3})$ complexity, such as matrix inversion, determinant calculation, or joint covariance estimation. We define the total execution time $T(d)$ as:
$T(d)=t_{ref}\cdot\left(\frac{d}{d_{ref}}\right)^{3}\cdot I,$ (21)
where $d_{ref}$ is the calibration dimension, $t_{ref}$ is the measured time per fundamental operation at $d_{ref}$, and $I$ is a conservative iteration constant (set to $100$ for minimum convergence).
C.2 Calibration Parameters
We performed the calibration on the same hardware used for CSB (NVIDIA RTX 3090). At $d_{ref}=50$, the measured time for a core $O(d^{3})$ operation (matrix inversion) was $t_{ref}\approx 0.000251$ seconds. Extrapolating to $d=10^{5}$:
$T(10^{5})$ $=0.000251\cdot\left(\frac{10^{5}}{50}\right)^{3}\cdot 100$ (22)
$=0.000251\cdot(2000)^{3}\cdot 100\approx 200,800,000\text{ seconds.}$ (23)
Converting to years: $200,800,000/31,536,000\approx\mathbf{6.37\text{ Years}}$.
C.3 The Memory Barrier (OOM)
Beyond time complexity, traditional methods face an immediate Memory Wall. A dense $10^{5}\times 10^{5}$ floating-point adjacency matrix requires $\approx 40$ GB of memory for storage alone. Standard algorithms requiring Hessian or Jacobian matrices would necessitate >400>400 GB, resulting in immediate Out-of-Memory (OOM) errors on current high-end GPUs. CSB bypasses this by operating solely on local factorized gradients, maintaining a constant memory footprint relative to node density.